# Framing Unionization on Facebook: Communication around Representation Elections in the United States

### v1

@@ Results
We applied our fine-tuned RoBERTa model to Facebook posts to perform multi-label classification of discourse frames. Table 1 reports performance in terms of precision, recall, and F1 scores for the test set. Scores are comparable to similar studies (Fernández-Zubieta et al. 2023; Mendelsohn et al. 2024), with F1 values ranging from 0.68 to 0.78 across frames and a macro F1 of 0.73.
@@
Table 1: Performance of the multi-label discourse frames classifier.
@@
Aiming at defining a baseline distribution of frame usage across unions, we first looked at the distributions of posts by union. We identified skewness (minimum = 779, 25th percentile = 1,921, 75th percentile = 5,442, maximum = 10,452, full distribution in Figure D1) and thus implemented a bootstrap-like down-sampling procedure. Over-represented unions were down-sampled to the minimum number of posts in the distribution, and the procedure was repeated with five different random seeds. For each frame, we computed the median proportion of posts expressing that frame across seeds. Overall, the median baseline distribution of frame usage is as follows: diagnostic in 40.9% of posts, community in 39.4%, prognostic in 28.4%, engagement in 23.1%, and motivational in 15.5%.
Figure 1 shows how individual unions diverge from the overall median baseline in their use of different frames. For each union, we calculated the share of posts expressing a given frame relative to its total output, and then expressed this value as a percentage difference over the median baseline. A value of +100% indicates that the union used that frame twice as often as the median baseline, 0% indicates equal usage, and –50% indicates that the union used the frame half as often. Building on a common distinction in labor movement scholarship (Herrigel 1993; Dubofsky 2000), we classified unions into two categories: craft unions, which organize workers sharing a specific trade or skill, and industrial unions, which organize workers across diverse trades, skill levels, and occupations. This classification was determined through manual review of information on unions’ official websites.
@@
Many unions behaved in line with the baseline; nonetheless, a clear divide emerged. Unions clustered into two broad groups based on whether their use of diagnostic and prognostic frames was higher or lower than the baseline. In particular, industrial unions tended to employ diagnostic and prognostic frames more frequently than the baseline. Examples include the Service Employees International Union (SEIU), the International Federation of Professional and Technical Engineers (IFPTE), the International Brotherhood of Teamsters (IBT), the Amalgamated Transit Union (ATU), National Nurses United (NNU), and Unite Here (UNITE HERE), all located toward the bottom of the dendrogram (union acronyms are expanded in Table D1). By contrast, lower-than-baseline usage of diagnostic and prognostic frames was more characteristic of craft unions. This distinction might reflect different organizational orientations: industrial unions, with their broad membership bases, are more often engaged in large-scale mobilization and bargaining arenas, where diagnostic frames (grievance articulation) and prognostic frames (solution proposals) play a central role (Voss and Sherman 2000). Variation also emerged in the use of the engagement frame. Industrial unions, in particular, often used engagement frames less than the baseline (cf. bottom of the plot), whereas higher-than-baseline engagement showed no consistent relationship to union structure (cf. top of the plot). Notably, the United Union of Roofers, Waterproofers and Allied Workers (Roofers) employed the engagement frame nearly two and a half times more often than the baseline.
@@
We next examined the usage of frames in the lead-up to election events. Figure 2 shows the average proportion of posts that express each frame in the pre-election period and its 95% confidence interval, separately for losing and winning outcomes. The usage of frames has been detrended by considering baseline periods with no election outcomes of a given kind: positive values indicate that a frame was used more often than usual, and negative values indicate lower-than-typical usage. This allowed us to examine how unions changed their frame usage leading up to elections relative to normal periods, and to compare won and lost cases. We down-sampled the number of wins and losses per union to ensure a balanced sample. Results are robust to the choice of sampling seed, as shown in Table C1.
Figure 2: Usage of frames leading to the election, comparing lost and won cases, in terms of mean and 95% confidence interval. Values on the y-axis refer to detrended scores. The $\star$ sign refers to the significance of the test for difference in means between losing and winning: $\star$ indicates significance at $\alpha=0.05$, $\star\star$ at $\alpha=0.01$, and $\star\star\star$ at $\alpha=0.001$.
In both won and lost cases, frame usage shifted before elections, but in different ways. Prior to wins, communication showed increases in all frames relative to baseline, except for engagement, which decreased. Prior to losses, by contrast, usage of motivational and community frames remained close to baseline, while diagnostic, prognostic, and engagement frames slightly decreased.
Next, we compared won and lost representation cases directly. Independent t-tests for the difference between group means showed statistically significant differences for all frames ($\alpha=0.001$, indicated as $\star$$\star$$\star$} in Figure 2), with communication in won cases exhibiting higher usage of nearly all frames compared to lost cases, particularly diagnostic, prognostic, and community frames. Robustness checks using multiple random seeds and down-sampling of over-represented unions showed that the differences in means between losses and wins were consistent for these three frames, while differences for motivational and engagement frames were not always significant (see Appendix Table C2).
@@
We finally examined how frame usage changes from the pre- to the post-election period. We first verified that frame usage before and after elections was robust to the choice of window size for rolling aggregation (see Appendix C.1). For each frame, we classified changes from pre- to post-election into three patterns: decrease, stable, or increase, depending on whether post-election usage was lower, similar, or higher than pre-election usage. We then counted the number of union–case instances with won or lost outcomes assigned to each pattern, providing an overall measure of how prevalent each change pattern was for each frame. We down-sampled the number of wins and losses per union to ensure a balanced sample. Results are robust to the choice of sampling seed, as shown in Figure C3. Figure 3 shows the difference in pattern prevalence between losses and wins, computed as losing minus winning. Positive values indicate a pattern was more common among losses, negative values indicate it was more common among wins. For example, a value of -5% in the decrease pattern for the prognostic frame means that in won cases, frame usage declined 5% more often than in lost cases. We computed Newcombe’s confidence intervals at $\alpha=0.05$ for the difference between the two independent proportions and considered a difference significant if the interval excluded 0.
Figure 3: Changes in frame usage before and after elections, comparing lost and won cases. Bold values in the cells refer to the difference between proportion of pattern presence for a given frame between losses and wins. Raw counts of union-case instances assigned to each cell are reported for losses (L) and wins (W). The $\star$ sign indicates a significant difference between losing and winning at $\alpha=0.05$.
In won cases, frame usage more often decreased for prognostic and motivational frames after elections. In lost cases, by contrast, prognostic and engagement usage more often increased, while motivational usage remained stable.
Robustness checks across multiple random seeds and down-sampling of over-represented unions confirmed that the decreases in prognostic and motivational frames in won cases were consistent. Other patterns were not consistently significant (see Appendix C.2).
@@
Second, in the lead-up to elections, unions in won cases systematically used nearly all frames more frequently than in lost cases, especially diagnostic, prognostic, and community (RQ2). This finding matches the literature on innovative strategies and union revitalization (Turner 2005), suggesting that successful unions may engage in more versatile and intensive framing strategies. It also aligns with research linking injustice frames and solidarity appeals to mobilization outcomes (Benford and Snow 2000; Pasquier and Wood 2018).
@@
Taken together, these findings connect directly to longstanding debates about revitalization of labor unions. Scholars have argued that unions must adapt not only organizationally but also communicatively (Milkman 2013), experimenting with innovative repertoires (Juravich and Bronfenbrenner 1998; Voss and Sherman 2000). By systematically distinguishing between the communication practices around successful and unsuccessful election outcomes, our study suggests how strategic framing might be an important part of these adaptation processes. More broadly, we show the value of micro-level, comparative framing analysis for advancing our understanding of how movements strategize and evolve in the digital era.
@@
While our study makes both methodological and substantive contributions, several limitations should be noted. First, our data is correlational, so observed associations between features and success cannot be interpreted causally. Second, the analysis is limited to Facebook—an important but partial element of unions’ communication strategies, which also involve other digital platforms and offline channels. Third, we focused on national and international level union accounts to ensure sufficient activity, which may obscure dynamics in smaller or more localized organizations. Fourth, we analyzed five discourse frames identified in prior work on social movements and civil society organizations. This enhances comparability but may overlook additional frames or nuances. Fifth, our pre- and post-election windows captured short-term dynamics but may miss longer-term adjustments. Finally, unions differ in size, resources, and political context, factors likely shaping communication in ways not fully accounted for here. We also acknowledge that systematic analyses of union communication could be misused by employers or political opponents, underscoring the need for ethical reflection in applying these methods.
@@
Despite these limitations, the study has implications for multiple audiences. For labor unions, the findings can inform digital communication strategies, highlighting how diagnostic, prognostic, community, motivational, and engagement frames can be used in election campaigns. For political communication researchers, the study offers large-scale empirical evidence on how civil society organizations employ framing online. For organizations more broadly, our results suggest ways to refine messaging to balance mobilization, solidarity, and engagement objectives. Methodologically, the released classifier and dataset provide a scalable resource for studying digital discourse framing across movements.
@@
Future work could extend this study in several directions. Expanding the frame taxonomy would allow more fine-grained analyses of communication strategies, including issue-specific frames. Examining union accounts at smaller, local levels—beyond the national and international accounts analyzed here—could reveal additional variation in messaging. Multi-platform analyses would provide a more holistic picture of unions’ communication ecosystems beyond Facebook. Incorporating a perspective on the targets of unions’ communication would also provide a broader understanding of strategic efforts, and complement existing literature suggesting ad hoc targeting as one valuable revitalization strategy (Earl 2019; Uba and Jansson 2021). Finally, incorporating audience responses through engagement metrics and comments would help link message production with resonance, offering deeper insight into the interplay between framing practices and collective action dynamics.
@@
With the large-scale adoption of digital technologies, labor unions have turned to online platforms as tools to engage with members, recruit new participants, and communicate with stakeholders. Scholars emphasize that unions’ strategies are shaped not only by external political contexts, but also by internal organizational choices (Martínez Lucio 2003). In this regard, Frege and Kelly (2004) identifies six strategic choices related to union revitalization, among which organizing—recruiting, representing, and retaining members, as well as expanding mobilization capacity—is particularly relevant in the digital context. Within these revitalization strategies, communication is central: discourse frames provide unions with a way to articulate grievances, propose solutions, and mobilize collective identities, thereby directly supporting organizing efforts. Building on this perspective, a consistent body of literature has linked framing and the labor movement.  Kelly (1998) adapts the collective action frames introduced by Snow, Benford et al. (1988) to the labor context and argues that successful organizing depends on workers interpreting grievances as unjust, identifying solutions, and believing collective action can be effective. Similarly, Voss and Sherman (2000) highlights the importance of identifying the right strategic repertoires of communication and engagement for the labor revitalization efforts to be successful. Other foundational work that followed highlights the importance of identifying the right frames for women and minorities (Bronfenbrenner 2005) and millenials (Milkman 2013).
@@
Despite these insights, evidence suggests that unions are not fully exploiting the potential of digital media.  Hodder and Houghton (2020) finds that online campaigning and news reporting remain dominant uses, leaving recruitment opportunities largely untapped. Likewise, Carneiro and Costa (2022) documents outdated strategies centered on one-way information sharing, rather than fostering interactive or targeted engagement. More promising results emerge when unions tailor their communication to specific worker groups (Jansson and Uba 2019) and account for the diversity of audiences with distinct communication needs (Panagiotopoulos 2021). These targeted strategies appear to be a necessary step towards building more effective digital communication.
@@
One line of work emphasizes the role of mobilization frames in online environments. Building on framing theory, scholars show how social media enable the rapid diffusion of diagnostic, prognostic, and motivational frames, which are crucial for lowering barriers to participation (Benford and Snow 2000; Fernández-Zubieta et al. 2023; Mendelsohn et al. 2024; Pera and Aiello 2025). A second perspective highlights the shift from the logic of collective action toward the logic of connective action (Bennett and Segerberg 2012). While traditional mobilization relies on formal organizations and collective identities, connective action underscores personalized, digitally mediated participation that can occur without strong organizational structures. This framework has been widely applied to cases such as the Arab Spring (Steinert-Threlkeld 2017) and climate activism (Segerberg and Bennett 2011). Finally, research on digital affordances investigates how platform features—such as sharing content, meta-voicing, and multimedia storytelling—shape the opportunities and constraints of collective action (Zheng and Yu 2016; Sæbø, Federici, and Braccini 2020).
@@
Acknowledgments
We acknowledge the support from the Carlsberg Foundation through the COCOONS project (CF21-0432). The funder had no role in study design, data collection and analysis, decision to publish, or preparation of the manuscript.

### v2

@@ Results
We applied our fine-tuned RoBERTa model to Facebook posts to perform multi-label classification of discourse frames. Table 1 reports performance in terms of precision, recall, and F1 scores for the test set. Scores are comparable to similar studies (Fernández-Zubieta et al. 2023; Mendelsohn et al. 2024), with F1 values ranging from 0.68 to 0.78 across frames and a macro F1 of 0.73. To evaluate the performance of this model, we compared it against two alternative classifiers, and found that it performs on par with or better than both (see Appendix D.1).
@@
Table 1: Performance of the multi-label discourse frames classifier. Support indicates the number of examples in each class or overall.
@@
Aiming at defining a baseline distribution of frame usage across unions, we first looked at the distributions of posts by union. We identified skewness (minimum = 779, 25th percentile = 1,921, 75th percentile = 5,442, maximum = 10,452, full distribution in Figure E1) and thus implemented a bootstrap-like down-sampling procedure. Over-represented unions were down-sampled to the minimum number of posts in the distribution, and the procedure was repeated with five different random seeds. For each frame, we computed the median proportion of posts expressing that frame across seeds. Overall, the median baseline distribution of frame usage is as follows: diagnostic in 40.9% of posts, community in 39.4%, prognostic in 28.4%, engagement in 23.1%, and motivational in 15.5%.
Figure 1 shows how individual unions diverge from the overall median baseline in their use of different frames. For each union, we calculated the share of posts expressing a given frame relative to its total output, and then expressed this value as a percentage difference over the median baseline. A value of +100% indicates that the union used that frame twice as often as the median baseline, 0% indicates equal usage, and –50% indicates that the union used the frame half as often.
@@
Many unions behaved in line with the baseline; nonetheless, a clear divide emerged. Unions clustered into two broad groups based on whether their use of diagnostic and prognostic frames was higher or lower than the baseline. In particular, industrial unions tended to employ diagnostic and prognostic frames more frequently than the baseline. Examples include the Service Employees International Union (SEIU), the International Federation of Professional and Technical Engineers (IFPTE), the International Brotherhood of Teamsters (IBT), the Amalgamated Transit Union (ATU), National Nurses United (NNU), and Unite Here (UNITE HERE), all located toward the bottom of the dendrogram (union acronyms are expanded in Table E1). By contrast, lower-than-baseline usage of diagnostic and prognostic frames was more characteristic of craft unions. This distinction might reflect different organizational orientations: industrial unions, with their broad membership bases, are more often engaged in large-scale mobilization and bargaining arenas, where diagnostic frames (grievance articulation) and prognostic frames (solution proposals) play a central role (Voss and Sherman 2000). Variation also emerged in the use of the engagement frame. Industrial unions, in particular, often used engagement frames less than the baseline (cf. bottom of the plot), whereas higher-than-baseline engagement showed no consistent relationship to union structure (cf. top of the plot). Notably, the United Union of Roofers, Waterproofers and Allied Workers (Roofers) employed the engagement frame nearly two and a half times more often than the baseline.
@@
Emotional Tone
To explore the affective nuance of frames, we computed correlations between frames and the emotional profiles of posts (cf. Figure 2).
Figure 2: Pearson’s correlation between frames and emotions in Facebook posts. Positive values (red) indicate positive correlations; negative values (blue) indicate negative correlations.
Diagnostic and prognostic frames were associated with stronger negative emotions, particularly anger and fear, while community frames were moderately correlated with joy. Motivational and engagement frames showed little association with emotions. Overall, correlations were modest, suggesting that emotions provide complementary context rather than replace frames as the primary lens for interpreting union communication strategies.
Topics Discussed
We also explored the thematic content of posts using topic modeling. Applying BERTopic to the full dataset yielded 308 topics with at least 50 posts each. Table 2 reports the ten most prevalent topics (excluding noise).
Topic % documents
Education and experience of teachers 3.81
Unions as communities 2.86
Updates and news from unions 2.47
Nurses and their importance 2.04
Congratulations to union members 1.95
Teamsters events 1.31
US postal service campaigns 1.27
Promotion of apprenticeships 1.03
Promotion of voting during US elections 0.99
SAG-AFTRA events 0.96
Table 2: Topic proportions for selected topics, considering all Facebook posts.
Some topics, such as unions as communities and promotion of apprenticeships, are relatively broad and span multiple unions. Others, like education and experience of teachers or nurses and their importance, are highly union-specific. At the union level, many topics remain highly specific (see Appendix C). These observations indicate that while topic modeling captures meaningful distinctions, prominent themes often reflect the particular context of individual unions. As a result, topic modeling is more useful for qualitative insight than for generalizable inference, and is therefore not included in the predictive model.
@@
We examined the use of frames in the lead-up to elections and their association with election outcomes using Bayesian logistic regression. To ensure balance, we down-sampled wins and losses within each union and fit the model across ten random seeds. To reduce collinearity, diagnostic and prognostic frame proportions were combined into a single category, as their Variance Inflation Factor (VIF) exceeded 5 when coded separately. Figure 3 presents posterior mean log-odds estimates for covariates whose posterior distributions excluded zero in at least six out of ten seeds.
Results are generally consistent across alternative modeling choices, including separate coding of diagnostic and prognostic frames, a different moving aggregation window size, and downsampling of over-represented unions (Figures D1D2D3, and D4).
Figure 3: Posterior mean log-odds estimates for selected variables, averaged across ten random seeds. Black dots indicate posterior means; grey lines show 90% Highest Density Intervals (HDIs) for each seed. Ratios denote the number of seeds (out of ten) with HDIs excluding zero. Variables shown meet this criterion in at least six seeds.
Greater use of diagnostic/prognostic and community frames prior to elections is associated with higher odds of winning, whereas motivational and engagement frames show no consistent association with the outcome. This pattern suggests that successful campaigns tend to emphasize problem definition, solution-oriented messaging, and collective identity rather than mobilization or engagement cues.
With respect to emotional content, neutral language is positively associated with winning, while disgust-related content is negatively associated; however, these effects are less stable across robustness checks than framing effects.
State-level random intercepts indicate heterogeneity in baseline odds of winning. Posterior means for Colorado, Idaho, and Ohio are shifted toward lower baseline odds relative to the global intercept, while Nevada, Rhode Island, and Washington exhibit higher baseline odds. The heterogeneity in state-level baseline odds may reflect differences in union density and organizing climate, with states such as Washington and Rhode Island exhibiting stronger labor infrastructures, while lower-density states may pose higher structural barriers to union success.
@@
We finally examined how frame usage changes from the pre- to the post-election period. We first verified that frame usage before and after elections was robust to the choice of aggregation window size (see Appendix D.3). For each frame, we classified changes from pre- to post-election into three patterns: decrease, stable, or increase, depending on whether post-election usage was lower, similar, or higher than pre-election usage. We then counted the number of union–case instances with won or lost outcomes assigned to each pattern, providing an overall measure of how prevalent each change pattern was for each frame. We down-sampled the number of wins and losses per union to ensure a balanced sample. Results are robust to the choice of sampling seed, as shown in Figure D5. Figure 4 shows the difference in pattern prevalence between losses and wins, computed as losing minus winning. Positive values indicate a pattern was more common among losses, negative values indicate it was more common among wins. For example, a value of -5% in the decrease pattern for the prognostic frame means that in won cases, frame usage declined 5% more often than in lost cases. We computed Newcombe’s confidence intervals at $\alpha=0.05$ for the difference between the two independent proportions and considered a difference significant if the interval excluded 0.
Figure 4: Changes in frame usage before and after elections, comparing lost and won cases. Bold values in the cells refer to the difference between proportion of pattern presence for a given frame between losses and wins. Raw counts of union-case instances assigned to each cell are reported for losses (L) and wins (W). The $\star$ sign indicates a significant difference between losing and winning ($p<0.05$).
In won cases, frame usage more often decreased for prognostic and motivational frames after elections. In lost cases, by contrast, prognostic and engagement usage increased more often, while motivational usage remained stable.
Robustness checks across multiple random seeds and down-sampling of over-represented unions confirmed that the decreases in prognostic and motivational frames in won cases were consistent. Other patterns were not consistently significant (see Appendix D.4).
@@
Second, we found that a higher usage of diagnostic, prognostic, and community frames in Facebook communication prior to elections was associated with higher odds of winning (RQ2). This finding matches the literature on innovative strategies and union revitalization (Turner 2005), suggesting that successful unions may engage in versatile and intensive framing strategies. It also aligns with research linking frames of injustice and solidarity to mobilization outcomes (Benford and Snow 2000; Pasquier and Wood 2018).
@@
Together, these findings connect directly to longstanding debates about revitalization of labor unions. Scholars have argued that unions must adapt not only organizationally but also communicatively (Milkman 2013), experimenting with innovative repertoires (Juravich and Bronfenbrenner 1998; Voss and Sherman 2000). By systematically distinguishing between the communication practices around successful and unsuccessful election outcomes, our study suggests that strategic framing may be an important part of these adaptation processes. More broadly, we show the value of micro-level, comparative framing analysis for advancing our understanding of how movements strategize and evolve in the digital era. It is important to note, however, that these conclusions are drawn from observational, correlational analyses and predictive modeling; as such, they do not support causal claims about the effects of specific framing strategies on election outcomes.
@@
While our study makes both methodological and substantive contributions, several limitations should be noted. First, our data is observational and correlational; consequently, observed associations between communication features and electoral success cannot be interpreted causally. In particular, unobserved organizational characteristics or strategic anticipation of election outcomes may jointly influence both communication practices and results. Second, the analysis is limited to Facebook—an important but partial element of unions’ communication strategies, which also involve other digital platforms and offline channels. Out of over 500 unions with available election outcomes, only 66 had Facebook accounts, and only 40 provided downloadable data, resulting in a relatively small and potentially skewed sample. Consequently, our findings may not generalize to all unions, particularly those for which Facebook is not a primary communication channel. Third, our analysis focuses on posts from the week immediately before and after each election. This window captures communication closely tied to electoral events and avoids noise from overlapping elections for a given union. However, it may miss broader or longer-term communication strategies, so results should be interpreted with this temporal limitation in mind. Fourth, we focused on national and international level union accounts to ensure sufficient activity, which may obscure dynamics in smaller or more localized organizations. Changes in data access policies could allow for analyses that can estimate the impact of such issues. Fifth, we analyzed five discourse frames identified in prior work on social movements and civil society organizations. This enhances comparability but may overlook additional frames or nuances. Finally, unions differ in size, resources, and political context, factors likely shaping communication in ways not fully accounted for here. We also acknowledge that systematic analyses of union communication could be misused by political actors, employers, or other organizations, underscoring the need for careful ethical reflection when applying these methods.
@@
Despite these limitations, the study has implications for multiple audiences. For labor unions, the findings can inform digital communication strategies, highlighting how different frames can be used in election campaigns. For political communication researchers, the study offers large-scale empirical evidence on how civil society organizations employ framing online. For organizations more broadly, our results suggest ways to refine messaging to balance mobilization, solidarity, and engagement objectives. Methodologically, the released classifier and dataset provide a scalable resource for studying digital discourse framing across movements.
@@
Future work could extend this study in several directions. Expanding the frame taxonomy would allow more fine-grained analyses of communication strategies, including issue-specific frames. Examining union accounts at smaller, local levels—beyond the national and international accounts analyzed here—could reveal additional variation in messaging. Multi-platform analyses would provide a more holistic picture of unions’ communication ecosystems beyond Facebook. Incorporating a perspective on the targets of unions’ communication would also provide a broader understanding of strategic efforts and complement existing literature that suggests ad hoc targeting as one valuable revitalization strategy (Earl 2019; Uba and Jansson 2021). Finally, incorporating audience responses through engagement metrics and comments would help link message production with resonance, offering deeper insight into the interplay between framing practices and collective action dynamics.
@@
With the large-scale adoption of digital technologies, labor unions have turned to online platforms as tools to engage with members, recruit new participants, and communicate with stakeholders. Scholars emphasize that unions’ strategies are shaped not only by external political contexts, but also by internal organizational choices (Martínez Lucio 2003). In this regard, Frege and Kelly (2004) identifies six strategic choices related to union revitalization, among which organizing—recruiting, representing, and retaining members, as well as expanding mobilization capacity—is particularly relevant in the digital context. Within these revitalization strategies, communication is central: discourse frames provide unions with a way to articulate grievances, propose solutions, and mobilize collective identities, thereby directly supporting organizing efforts. Building on this perspective, a consistent body of literature has linked framing and the labor movement.  Kelly (1998) adapts the collective action frames introduced by Snow, Benford et al. (1988) to the labor context and argues that successful organizing depends on workers interpreting grievances as unjust, identifying solutions, and believing collective action can be effective. Similarly, Voss and Sherman (2000) highlights the importance of identifying the right strategic repertoires of communication and engagement for the labor revitalization efforts to be successful. Other foundational work that followed highlights the importance of identifying the right frames for women and minorities (Bronfenbrenner 2005), and millenials (Milkman 2013).
@@
Despite these insights, evidence suggests that unions are not fully exploiting the potential of digital media.  Hodder and Houghton (2020) finds that online campaigning and news reporting remain dominant uses, leaving recruitment opportunities largely untapped. Likewise, Carneiro and Costa (2022) documents outdated strategies centered on one-way information sharing, rather than fostering engagement. More promising results emerge when unions tailor their communication to specific worker groups (Jansson and Uba 2019) and account for the diversity of audiences with distinct communication needs (Panagiotopoulos 2021). These targeted strategies appear to be a necessary step towards building more effective digital communication.
@@
One line of work emphasizes the role of mobilization frames in online environments. Building on framing theory, scholars show how social media enables the diffusion of diagnostic, prognostic, and motivational frames, crucial for lowering barriers to participation (Benford and Snow 2000; Fernández-Zubieta et al. 2023; Mendelsohn et al. 2024; Pera and Aiello 2025). A second perspective highlights the shift from the logic of collective action toward the logic of connective action (Bennett and Segerberg 2012). While traditional mobilization relies on formal organizations and collective identities, connective action underscores personalized, digitally mediated participation that can occur without strong organizational structures. This framework has been applied to cases such as the Arab Spring (Steinert-Threlkeld 2017) and climate activism (Segerberg and Bennett 2011). Finally, research on digital affordances investigates how platform features—such as sharing content, meta-voicing, and multimedia storytelling—shape the opportunities and constraints of collective action (Zheng and Yu 2016; Sæbø, Federici, and Braccini 2020).